Two buses leave the station at 4:00A.M, one travelling east at 60kph and
the other, west at 65kph. At what time will they be 50km apart? Please Include Represent, Relate, Equate, Solve and Prove. Thank you!

Answer by ankor@dixie-net.com(22740)

Two buses leave the station at 4:00A.M, one travelling east at 60kph and
the other, west at 65kph.
At what time will they be 50km apart?
:
Let t = no. of hrs they traveled to be 50 km apart
:
Write a dist equation, dist = speed * time
60t + 65t = 50
125t = 50
t = 50/125
t = .4 hrs
:
.4 hrs = .4(60) = 24 min
therefore
at 4:24 Am they will be 50 km apart
:
Check this, find the dist each traveled
.4 * 60 = 24 km
.4 * 65 = 26 km
------------------
total d: 50 km